OpenStudy (anonymous):

Which of these is a simplified form of the equation 7x + 4 = 5 + 3x + 1x? 3x = 9 3x = 1 11x = 9 11x = 1

Mehek (mehek14):

combine like terms \(7x+4=5+3x+1x\\3x+1x=?\)

Mehek (mehek14):

@yvyvvyv ?

OpenStudy (anonymous):

how do i do that

Mehek (mehek14):

3 + 1?

OpenStudy (anonymous):

4

Mehek (mehek14):

ok and then just add the variable so 4x \(7x + 4 = 5 + 4x\) 5 - 4?

OpenStudy (anonymous):

1

Mehek (mehek14):

7x - 4x (7 - 4)

OpenStudy (anonymous):

3

Mehek (mehek14):

add the variable \(3x = 1\)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

no 1

Mehek (mehek14):

5 - 4 = 1 7x - 4x = 3x \(3x = 1\)

OpenStudy (anonymous):

oh okay I got it